OpenStudy (anonymous):

If kx+y-2=0 and x+hy-3=0 represent the same line, find h and k.

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

To solve for h, Same line = have the same y-intercept. How to find the y-intercept from the given equation? To solve for k, Same line = have the same x-intercept. How to find the x-intercept from the given equation?

OpenStudy (anonymous):

ummm...

OpenStudy (callisto):

Given Ax+By+C =0, slope = -A/B x-intercept = -C/A y-intercept = -C/B

OpenStudy (anonymous):

i am just thinking of these formula...

OpenStudy (callisto):

Now, set up the equations and solve the unknowns

OpenStudy (anonymous):

wait a moment, let me calculate it

OpenStudy (callisto):

Take your time :)

OpenStudy (anonymous):

do we need the formula y=mx+c? for k, slope=-k y-intercept=2 x-intercept=2/k

OpenStudy (anonymous):

for h, slope=-1/h x-intercept=3 y-intercept=3/h

OpenStudy (anonymous):

oh, then 2=3/h h=3/2 and 2/k=3 k=2/3 right?

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

thank you so much!

OpenStudy (callisto):

You're welcome :)